In a regular hexagon ABCDEF, if \overrightarrow{AB}.=\vec{a} and vector BC = vector b,


then express vector CD, vectorDE, vectorEF, vectorFA, vectorAC, vectorAD, vectorAF, vectorCF IN TERMS OF \vec{a} AND VEC B ?

Grade:12

3 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Aby Alex

given AB =a

  BC =b      here a and b are vector

EF=-b

DE=-a

and we know thar AD=2BC

                             =2b

AD=AB+BC+CD

2b =a+b+CD

CD =b-a

and FA=-CD

           =a-b

Ac=AB+BC

    =a+b

AF=CD=b-a

and CF=2DE

          =-2a


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

 

Aby Alex
14 years ago

Thank u very much

Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to the problem.
 
Given,
AB = a
BC = b      here a and b are vector
EF = – b
DE = – a
and we know thar AD = 2BC
                                =2b
AD = AB+BC+CD
2b = a+b+CD
CD = b – a
and FA = – CD
           = a – b
AC = AB+BC
     = a+b
AF = CD
     = b – a
and CF = 2DE
            = – 2a
 
Thanks and regards,
Kushagra